OpenStudy (anonymous):

a loudspeaker has a radius and a height of 7.5 cm calculate the volume of the cone

OpenStudy (cwrw238):

volume of a cone = (1/3) pi r^2 h

OpenStudy (anonymous):

can u explain it properly please @cwrw238

OpenStudy (anonymous):

with workin out thanks

OpenStudy (cwrw238):

r = radius , h = height , use pi = 3.14

OpenStudy (jdoe0001):

\(\bf \large \textit{volume of a cone} = \cfrac{\boldsymbol{\pi} r^2 h}{3}\\ \quad \\ r = radius\\ h = height\)

OpenStudy (cwrw238):